OpenStudy (anonymous):

Factor the algebraic expression (x - 1)2 - (y - 2)2. WILL FAN AND MEDAL

OpenStudy (anonymous):

@TheSmartOne

rishavraj (rishavraj):

a^2 - b^2 = (a - b) (a + b)

OpenStudy (anonymous):

so what is a and b?

OpenStudy (anonymous):

or how do we find out?

rishavraj (rishavraj):

ii ur question a = x - 1 b = y - 2

OpenStudy (anonymous):

okay so now what?

rishavraj (rishavraj):